Consider the following equation:


Ca(OH)2(s) + 2 HCl(aq) —> CaCl2(aq) + 2 H2O(l)



a) How many liters of 0.100 M HCl is required to completely react with 5.00 grams of calcium hydroxide?



b) If 15.0 grams of calcium hydroxide is combined with 75.0 mL of 0.500 M HCl, how many grams of calcium chloride


would be formed?

1
Expert's answer
2022-03-22T00:56:04-0400

Moles of calcium hydroxide

Moles= 5/74= 0.0676

(0.0676/0.5)× 1000= 135ml